# SimPro: A Simple Probabilistic Framework Towards Realistic Long-Tailed Semi-Supervised Learning

### v2

@@ Appendix B More Experimental Results
Appendix C Proof of Proposition 1
@@
We employ the method of Lagrange multipliers to find the optimal values of $\bm{\hat{\pi}}_{l}$ and $\bm{\hat{\pi}}_{u}$ that maximize the $\mathcal{Q}$ function subject to the constraints of probability distributions (i.e., the elements of $\bm{\pi}_{l}$ and $\bm{\pi}_{u}$ must sum to 1). Let $\lambda_{l}$ and $\lambda_{u}$ be the Lagrange multipliers for these constraints. The Lagrangian $\mathcal{L}$ can be formulated as 111 $y$ is one-hot label, but when $y$ is the subscript of a variable, it represents the $y$-th category, that is, $\phi_{y}=\phi_{\mathop{\mathrm{argmax}}y}$. we omit this difference without affecting the understanding.:
@@
$-\lambda_{l}\left(\sum_{y}\pi_{ly}-1\right)-\lambda_{u}\left(\sum_{y}\pi_{uy}-1\right).$
@@
Appendix D Proof of Proposition 2
@@
Appendix E Proof of Proposition 3
@@
$P_{d}(y|x)$ $\propto\exp(f_{\bm{\theta}}(x,y))$ (49)
$\propto\frac{\phi^{*}_{y}}{\hat{\phi}_{y}}P(x|y)$ (50)
@@
It is evident that minimizing the decision error rate is achieved when $\bm{\hat{\phi}}=\bm{\phi}^{*}$, leading to the Bayes decision error rate, which is the theoretical lower bound of error rates across all possible distributions.
@@
$P(e;\bm{\hat{\phi}})-\inf_{\bm{\phi}}P(e;\bm{\phi})$ $=\int_{\lambda\leq l(x)\leq 1}|P(-1|x)-P(+1|x)|P(x)\,\mathop{}\!\mathrm{d}x$ (56)
$=\int_{\lambda\leq l(x)\leq 1}|1-l(x)|P(-1|x)P(x)\,\mathop{}\!\mathrm{d}x$ (57)
$\leq(1-\lambda)\int_{\lambda\leq l(x)\leq 1}P(-1,x)\,\mathop{}\!\mathrm{d}x$ (58)
$\leq(1-\lambda)\int_{x}P(-1,x)\,\mathop{}\!\mathrm{d}x$ (59)
$=(1-\lambda)P(-1)$ (60)
$=\frac{1-\lambda}{2}$ (61)
$=\frac{\hat{\phi}_{-1}-\phi^{*}_{-1}}{2\phi^{*}_{+1}\hat{\phi}_{-1}}$ (62)
$\leq\frac{1}{2\phi^{*}_{+1}\phi^{*}_{-1}}(\hat{\phi}_{-1}-\phi^{*}_{-1})$ (63)

### v3

@@ 4.2 Analysis
(a) Marginal distribution estimation. We first investigate the impact of the estimation for $P(y|\bm{\pi})$ (M-step, Prop. 1) and its usage in building the Bayes classifier for pseudo-label generation (E-step, Prop. 2). The results in Tab. 5 substantiate the high effectiveness and the necessity of this estimation in driving the success of our approach, thereby validating our methodology both theoretically and practically.
Hyperparameter Sensitivity.
As outlined in App. B, we discover that reducing the threshold value improves performance for the CIFAR100-LT dataset. The rationale behind adjusting the confidence threshold is based on the observation that an increase in the number of classes typically results in a corresponding decrease in the confidence of the predictions. Consequently, it becomes necessary to lower the threshold to accommodate this change in confidence levels. Sensitivity analysis regarding the threshold value is presented in Fig. 4. It is consistently observed across different settings that the optimal performance is achieved when the threshold is set at $t=0.2$ and $t=0.95$ for the CIFAR100-LT and CIFAR10-LT, respectively. Moreover, to compare with ACR, we also conduct a sensitivity analysis of the confidence threshold $t$ for ACR. The results in Fig. 6 of App. C demonstrate that lowering the threshold does not improve performance for ACR.
Visualization of Estimation Quality.
Our study includes a visualization of the estimated distribution quality in Fig. 5. The vertical axis quantifies the Kullback-Leibler (KL) divergence, which measures the deviation of the estimated distribution from the ground truth. The results indicate a significant improvement in estimation accuracy after only a few training epochs. This empirical evidence validates the effectiveness of the theoretically derived estimation method of distribution, as outlined in Prop. 1.
5 Conclusion
In this paper, we introduce SimPro, a novel probabilistic framework for realistic long-tailed semi-supervised learning (ReaLTSSL). This framework represents a pioneering advancement in the field by innovatively enhancing the Expectation-Maximization (EM) algorithm. The key innovation lies in the explicit separation of the estimation process for conditional and marginal distributions. In the M-step, this separation allows for the derivation of a closed-form solution for the marginal distribution parameters. Additionally, SimPro optimizes the parameters of the conditional distribution via gradient descent, facilitating the learning of a Bayes classifier. In the E-step. the Bayes classifier, in turn, generates high-quality pseudo-labels. SimPro is characterized by its elegant theoretical underpinnings and its ease of implementation, which requires only minimal modifications to existing codebases. Furthermore, we incorporate two innovative class distributions specifically for unlabeled data. These distributions provide a more comprehensive and realistic evaluation of ReaLTSSL algorithms. Empirical evidence from various benchmarks demonstrates that SimPro consistently delivers state-of-the-art performance, highlighting its robustness and superior generalization capabilities.
Acknowledgements
This work was supported in part by the National Key R&D Program of China under Grant 2021ZD0140407, in part by the National Natural Science Foundation of China under Grants 62276150, 62321005 and 42327901.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Details of the Probabilistic Model
We provide a detailed derivation and analysis to demonstrate that the probabilistic model is correctly defined with an explicit parameterization of $p(x;\xi)$. Given the independence of parameters through conditional parameterization (Koller & Friedman, 2009), we can decompose the joint probability distribution as follows:
$p(x,y;\theta,\pi,\xi)=p(x;\xi)p(y|x;\theta,\pi)=p(y;\pi)p(x|y;\theta,\xi).$ (18)
Applying Bayes’ rule, we obtain:
$\frac{p(y|x;\theta,\pi)}{p(y;\pi)}=\frac{p(x|y;\theta,\xi)}{p(x;\xi)}.$ (19)
It is evident that $\pi$ and $\xi$ appear only on the left and right sides of the equation, respectively, indicating that the equation is neither a function of $\pi$ nor $\xi$ but is parameterized solely by $\theta$. We define the above equation as $g(x,y;\theta)$, that is:
$\frac{p(x|y;\theta,\xi)}{p(x;\xi)}=\frac{p(y|x;\theta,\pi)}{p(y;\pi)}=g(x,y;\theta).$ (20)
Returning to the equation in the main paper, we have:
$p(y|x;\theta,\pi)=\frac{p(y;\pi)p(x|y;\theta,\xi)}{p(x;\xi)}=p(y;\pi)g(x,y;\theta).$ (21)
Although we explicitly parameterize $p(x;\xi)$, it is clear that $p(y|x;\theta,\pi)$ is parameterized solely by $\theta$ and $\pi$, and is independent of $\xi$. In fact, the fitting target of the network parameters $\theta$ is $g(x,y;\theta)=p(x|y;\theta,\xi)/p(x;\xi)$.
Since we did not explicitly parameterize $p(x)$ in our framework, when $x$ is not a condition, the parameters of the relevant notions omit the parameters of $p(x)$, such as $p(x)$, $p(x|y)$, $p(x,y)$, etc. This may lead to a potential misunderstanding, as:
$p(x)=\sum_{y}p(x|y;\theta)p(y;\pi)$ (22)
suggests that $p(x)$ seems to be parameterized by $\theta$ and $\pi$. However, if we recover the omitted parameter of $p(x)$, we have:
$p(x;\xi)=\sum_{y}p(x|y;\theta,\xi)p(y;\pi)=p(x;\xi)\sum_{y}g(x,y;\theta)p(y;\pi)=p(x;\xi)\sum_{y}p(y|x;\theta,\pi)=p(x;\xi),$ (23)
which is consistent with the explicit parameterization of $p(x;\xi)$.
Therefore, the probabilistic model is correctly defined without the explicit parameterization of $p(x)$.
Appendix B Experimental Setup
Datasets.
To validate the effectiveness of SimPro, we employ five commonly used SSL datasets, CIFAR10/100 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), ImageNet-127 (Fan et al., 2022) and orginal ImageNet-1k (Deng et al., 2009). Following the methodology described in ACR (Wei & Gan, 2023), we denote the number of samples for each category in the labeled, unlabeled dataset as $N_{1}\!\geq\!\cdots\!\geq\!N_{K}$, $M_{1}\!\geq\!\cdots\!\geq\!M_{K}$, respectively, where $1,\cdots,K$ are the class indices. We define $\gamma_{l}$, $\gamma_{u}$, and $\gamma_{t}$ as the class imbalance ratios for labeled, unlabeled, and test data, respectively. We specify ‘LT’ for those imbalanced variants. These ratios are calculated as $\gamma_{l}\!=\!N_{1}/N_{K}$ and $\gamma_{u}\!=\!M_{1}/M_{K}$. The sample number of the $k$-th class follows an exponential distribution, expressed as $N_{k}\!=\!N_{1}\cdot\gamma_{l}^{-\frac{k-1}{K-1}}$ for labeled and $M_{k}\!=\!M_{1}\cdot\gamma_{u}^{-\frac{k-1}{K-1}}$ for unlabeled data.
As illustrated in Fig. 1, for the CIFAR10/100 datasets, we constructed five class distributions to test the performance of different algorithms under more general settings. Regarding the STL10 dataset, due to the unknown ground-truth labels of the unlabeled data, we approach the experiments by controlling the imbalance ratio of the labeled data.
For ImageNet-127, we follow the original setting in Fan et al. (2022) ($\gamma_{l}\!=\!\gamma_{u}\!=\!\gamma_{t}\approx 286$). Nevertheless, this approach does not serve as an appropriate benchmark for long-tail learning. In these scenarios, biased classifiers often exhibit high performance, which is exactly the outcome we seek to prevent. Consequently, we also resample the test dataset to ensure a uniform distribution ($\gamma_{t}\!=\!1$). Following Fan et al. (2022), we reduce the image resolution to $32\times 32$ and $64\times 64$ in response to resource constraints.
Training hyper-parameters.
Our experimental setup mainly follows FixMatch (Sohn et al., 2020) and ACR (Wei & Gan, 2023). For example, we employ the Wide ResNet-28-2 (Zagoruyko & Komodakis, 2016) on CIFAR10/100 and STL10, and ResNet-50 (He et al., 2016) on ImageNet-127. All models are optimized with SGD. Several settings are slightly different from those in ACR: as outlined in Sec. 3.3, to achieve a balanced training, we set the ratio between unlabeled and labeled samples in each batch as $\mu\!=\!M/N$ ($8$ on CIFAR, $16$ on STL10-LT, and $2$ on ImageNet-127, ), where $M,N$ are the total sample numbers of unlabeled/labeled data. In contrast, $\mu$ is set as $2$ for all datasets in ACR. The effectiveness of this adjustment is validated in Tab. 6.
The batch size for labeled data is $64$ on CIFAR10/100 and STL10-LT, and $32$ on ImageNet-127. To ensure a fair comparison, the training epochs are reduced to $86$ on CIFAR10/100 and STL10-LT, and $500$ on ImageNet-127. The initial learning rate $\eta$ is linearly scaled to $0.17$ on CIFAR10/100 and STL10-LT, and $0.01$ on ImageNet-127, which decays with a cosine schedule (Loshchilov & Hutter, 2017) as in ACR.
Regarding the hyperparameter $\tau$ used in Eq. 14, we follow the guidelines from Menon et al. (2021) and set $\tau\!=\!2.0$ and $1.0$ for CIFAR10-LT/STL10-LT and CIFAR100-LT/ImageNet-127, respectively.
For the confidence threshold $t$ in Eq. 1), we set $t\!=\!0.95$ on CIFAR10-LT/STL10-LT following Sohn et al. (2020). We adjust it to $0.2$ on CIFAR100-LT/Imagenet-127, as we observe that reducing the threshold enhances performance (Fig. 4).
Specifically, the settings on ImageNet-1k are identical to those on ImageNet-127.
Appendix C More Experimental Results
Table 8: The impact of the predefined anchor distribution in ACR (Wei & Gan, 2023) on CIFAR10-LT with $\gamma_{l}=100,N_{1}=500$, and $M_{1}=4000$. $\star$ denotes that we use the predefined anchor distributions to estimate $P(y|\bm{\pi})$ in our SimPro.
consistent uniform reversed middle head-tail
ACR 81.6 92.1 85.0 73.6 79.8
SimPro 80.7 93.8 85.8 84.8 83.0
SimPro 82.7 94.3 86.0
Table 9: The impact of the predefined anchor distribution in ACR (Wei & Gan, 2023) on CIFAR100-LT with $\gamma_{l}=20,N_{1}=50$, and $M_{1}=400$. $\star$ denotes that we use the predefined anchor distributions to estimate $P(y|\bm{\pi})$ in our SimPro.
consistent uniform reversed middle head-tail
ACR 44.9 52.2 42.3 42.6 42.6
SimPro 43.1 52.3 45.5 43.6 44.8
SimPro 45.9 53.8 46.0
Figure 6: Sensitive analysis of the confidence threshold $t$ for ACR (Wei & Gan, 2023) on CIFAR100-LT with $\gamma_{l}=20$, $N_{1}\!=\!50$, and $M_{1}\!=\!400$ and CIFAR10-LT with $\gamma_{l}\!=\!150$, $N_{1}\!=\!500$, and $M_{1}\!=\!4000$. The optimal performance is achieved across different settings when the threshold is set at $0.95$.
Appendix D Proof of Proposition 1
@@
[…]
@@
$-\lambda_{l}\left(\sum_{y}\pi_{ly}-1\right)-\lambda_{u}\left(\sum_{y}\pi_{uy}-1\right).$ (24)
@@
Appendix E Proof of Proposition 2
@@
Appendix F Proof of Proposition 3
@@
$P_{d}(y|x)$ $\propto\exp(f_{\bm{\theta}}(x,y))$
$\propto\frac{\phi^{*}_{y}}{\hat{\phi}_{y}}P(x|y)$
@@
Minimizing the decision error rate is achieved when $\bm{\hat{\phi}}=\bm{\phi}^{*}$, leading to the Bayes decision error rate, which is the theoretical lower bound of error rates across all possible distributions.
@@
$P(e;\bm{\hat{\phi}})-\inf_{\bm{\phi}}P(e;\bm{\phi})$ $=\int_{\lambda\leq l(x)\leq 1}|P(-1|x)-P(+1|x)|P(x)\,\mathop{}\!\mathrm{d}x$
$=\int_{\lambda\leq l(x)\leq 1}|1-l(x)|P(-1|x)P(x)\,\mathop{}\!\mathrm{d}x$
$\leq(1-\lambda)\int_{\lambda\leq l(x)\leq 1}P(-1,x)\,\mathop{}\!\mathrm{d}x$
$\leq(1-\lambda)\int_{x}P(-1,x)\,\mathop{}\!\mathrm{d}x$
$=(1-\lambda)P(-1)$
$=\frac{1-\lambda}{2}$
$=\frac{\hat{\phi}_{-1}-\phi^{*}_{-1}}{2\phi^{*}_{+1}\hat{\phi}_{-1}}$
$\leq\frac{1}{2\phi^{*}_{+1}\phi^{*}_{-1}}(\hat{\phi}_{-1}-\phi^{*}_{-1})$